# Tree-of-Code: A Tree-Structured Exploring Framework for End-to-End Code Generation and Execution in Complex Task Handling

### v1

@@ 4.3 Final Result Generator
Once valid outputs from successfully executed nodes are collected, the same LLM makes the final decision by performing a majority vote and summarization to determine the most likely answer — this corresponds to the green node in Figure 3. Other valid responses (red nodes) are discarded.
@@
ReAct CodeAct CodeProgram ToC (1-3) ReAct CodeAct CodeProgram ToC (1-3)
@@
Table 1: Detailed performance comparison of different models under ReAct, CodeAct, CodeProgram, and one-layer, three-node Tree-of-Code (ToC) on the M3ToolEval and API-Bank level-3 datasets. The correctness is reported, with the average number of turns in parentheses. The noteworthy point is that since the model is fixed, the ToC mechanism at this time represents the ToC mechanism without model exploration (w/o model exploration).
5 Experiment and Analysis
5.1 Setup
Datasets. Following CodeAct, our evaluation is based on M3ToolEval [^1] https://github.com/xingyaoww/code-act/tree/main/scripts/eval/m3tooleval (M3) Wang et al. (2024) and the test set of API-Bank [^2] https://huggingface.co/datasets/liminghao1630/API-Bank/tree/main Li et al. (2023b). M3 consists of 82 tasks utilizing 100 tools in code/JSON/txt action space respectively across 5 types of scenarios, including DNA sequencer, message decoder, trade calculator, travel itinerary planning, and web browsing. API-Bank contains 314 tool-use dialogues and 73 API tools, including level-1, 2, 3. Unlike CodeAct, which evaluates only on level-1, we focus directly on the 50 most challenging level-3 tasks, on which nearly all non-GPT4 models score 0%, according to the original paper.
For M3, we add the "next_action" tool, which is a customized function introduced in Section 3.2. For API-Bank, which only supports JSON format, we make the following modifications to adapt it for code interaction: (1) functionalize all API tools, (2) add output examples to each function description (shown in Figure 7). We include all tool signatures in the prompt context and let LLMs inherently search and select tools, instead of using ToolSearch API, deemed the least essential in Li et al. (2023b). (3) determine correctness by matching the response to the expected final output through conditional keywords, not by API call matching.
Baselines. ReAct Yao et al. (2022) combines reasoning and action in a dynamic, step-by-step interaction, providing a flexible approach to task-solving by adjusting action strategies based on environmental feedback. CodeAct Wang et al. (2024) replaces the JSON in ReAct with a block of code as the LLM agent’s action, enabling multi-turn interactions and effectively expanding the action space for solving complex real-world problems.
Metrics. The evaluation includes accuracy and averaged turns. Accuracy represents the percentage of complex tasks that are correctly solved. Each LLM-generated code is considered one turn. For parallel generation, the number of threads counts as turns in terms of resource usage, but when considering time, multiple parallel generations count as one turn. We use the latter approach.
5.2 One-turn vs. Multi-turn
CodeProgram enables global planning and complete solutions in a single turn by leveraging code’s ability to handle long logic chains, aligning with global reasoning in language, and defining clear tool inputs and outputs. With a significant advantage in the number of turn, Table 1 and Figure 6 demonstrate that the performance of some models even surpasses the multi-turn CodeAct, particularly for the Claude series models. We grow CodeProgram into a single-layer, three-node (1-3) Tree-of-Code (ToC). The prompt for each node is randomly sampled from our prompt pool. Compared to CodeProgram, the simple 1-3 tree-structured ToC with random prompts significantly boosts performance. The average performance of 1-3 ToC already surpasses CodeAct, highlighting the power of prompt randomness. Upon reviewing the generated code, we observe that LLMs often produce modular code for each step or add comments before modules, even when not explicitly required.
Figure 7: Example of the function signature in level-3.
5.3 ToC vs. CodeAct and ReAct
We primarily compare the ToC framework, which is comprised of CodeProgram nodes, with the CodeAct and ReAct framework, which are comprised of steps, using the M3 and the level-3 datasets. For ToC, we randomly sample the LLM and prompt from the LLM list and prompt pool, respectively, at each node exploration. For CodeAct and ReAct, we report the average results across all models used in this paper. Table 2 shows that ToC consistently achieves superior performance and demonstrates a significant advantage with fewer interaction steps, highlighting its efficiency in managing complex tool-use scenarios.
Mechanism M3ToolEval API-Bank level-3
Avg Turns Correct Avg Turns Correct
ReAct 8.2 38.1% 9.5 8.2%
CodeAct 7.0 49.4% 8.9 19.2%
Tree-of-Code 1.7 67.1$\uparrow$ 2.1 38.0% $\uparrow$
Table 2: Performance comparison of CodeAct and our ToC in terms of averaged turns and accuracy on M3 and API-Bank level-3 tasks. Note: all numerical results presented in this paper are rounded to one decimal place.
5.4 Analysis and Ablation Studies
@@
Visualization results for ReAct, CodeAct, and 3-3 ToC on the M3 dataset (Figure 8) show that ToC achieves near-perfect accuracy on all tasks except the web browsing task.
Figure 8: Comparison across five tasks in the M3.
Prompt exploration. The ablation results in Table 4 demonstrate the effectiveness of prompt exploration. By comparing the random model with the fixed model, prompt exploration proves more crucial in low-diversity scenarios.
@@
Table 4: Results of the ablation study.
6 Discussion
Why we do not try a search tree? We initially explored using evolutionary algorithms to merge nodes from different branches for the next generation, aiming to reduce the search space. However, after extensive testing, we found this approach impractical and hard to implement. We analyzed error cases from ToC and found that a single branch can easily fall into specific errors. Since different branches follow distinct reasoning paths, they struggle to learn from each other. Even with reflection and random exploration, LLMs are prone to early-stage errors that disrupt subsequent reasoning An et al. (2023); Bao et al. (2024); Tong et al. (2024). This issue mirrors the human cognitive challenge of breaking out of a "bistable state" — like the famous "duck-rabbit illusion" or Rubin’s vase Hancock (2013). In such states, a person may generate multiple parallel thoughts, but once one is chosen, it is difficult to switch to another without external intervention Andreev et al. (2020). This highlights the need for introducing random exploration and multiple nodes at each growth step, as these strategies help overcome cognitive bottlenecks.
@@ 7 Related Work
LLM Code Generation. Chain of Codes framework Li et al. (2023a) expands the range of reasoning tasks that LLMs can solve by "thinking in code." Similarly, CodePlan Wen et al. (2024a) utilizes pseudocode to guide structured reasoning. Additionally, the Structured Chain-of-Thought Prompting (SCoT) technique Esfahani et al. (2024) has highlighted the potential of structured prompts in this domain. Recent works combining code with agents have primarily focused on task completion within programming-related domains, such as software development Qian et al. (2024); Wang et al. (2023), programming assistance Islam et al. (2024); Wen et al. (2024c), and scientific problems Chen et al. (2022); Gao et al. (2023); Hong et al. (2024). Few methods Wang et al. (2024) treat code as a scalable language format to call multiple tools for solving complex real-world tasks.
Recently, we found a new work, CodeTree Li et al. (2024b), which uses a tree structure to explore the search space of code generation tasks. Unlike our approach, it focuses on multi-agent searching rather than an end-to-end self-growing tree. Additionally, it was released three months later than the initial submission of our work.
@@ 8 Conclusion
In this paper, we introduced the Tree-of-Code (ToC) method, which combines execution-based reflection with end-to-end thought-code generation for complex task handling. With efficient model integration and prompt exploration, ToC significantly outperformed the baselines on two complex task datasets, boosting both efficiency and task-solving capabilities. The ToC framework opens up exciting possibilities for advancing human-like global cognition, inspiring further exploration of tree structures in end-to-end code generation, particularly for complex multi-tool tasks, data generation, and reinforced fine-tuning.
@@
Additional engineering effort
We efficiently transform multi-turn interactions into a single-turn complete program comprising a series of tool calls. However, this approach demands more detailed usage instructions for the tools. For actions with limited semantic information (e.g., webpage clicking and scrolling), an artificially constrained action exploration space is also necessary. Consequently, additional engineering efforts are required to fully implement our end-to-end code generation approach.
@@
We emphasize that our method operates at the granularity of code "program" rather than "action". However, it is limited in fully open-ended scenarios requiring step-by-step exploration, such as a robot navigating an unfamiliar environment, or in handling tasks with extremely long sequences beyond the capabilities of current reasoning methods, like generating an entire paper. In such cases, it cannot provide a complete final solution. For larger and more complex system programs in the future, our method may serve as a "subprogram" within the overall solution, similar to a single agent’s role in multi-agent systems.

### v2

@@ 3.4 Final Result Generator
Once valid outputs from successfully executed nodes are collected, the same LLM makes the final decision by performing a majority vote and summarization to determine the most likely answer. Ties are rare in our observations, so we always choose the most frequent answer without special handling.
4 Experiment and Analysis
4.1 Setup
Datasets. Following CodeAct, our evaluation is based on M3ToolEval [^1] https://github.com/xingyaoww/code-act/tree/main/scripts/eval/m3tooleval (M3) Wang et al. (2024) and the test set of API-Bank [^2] https://huggingface.co/datasets/liminghao1630/API-Bank/tree/main Li et al. (2023b). M3 consists of 82 tasks utilizing 100 tools in code/JSON/txt action space respectively across 5 types of scenarios, including DNA sequencer, message decoder, trade calculator, travel itinerary planning, and web browsing. API-Bank contains 314 tool-use dialogues and 73 API tools, including level-1, 2, 3. Unlike CodeAct, which evaluates only on level-1, we focus directly on the 50 most challenging level-3 tasks, on which nearly all non-GPT4 models score 0%, according to the original paper. Considering API-Bank only supports JSON format, we make following modifications to adapt it for code interaction: (1) functionalize all API tools, (2) add output examples to each function description (Figure 6). We include all tool signatures in the prompt context and let LLMs inherently search and select tools, instead of using ToolSearch API, deemed the least essential in Li et al. (2023b). (3) determine correctness by matching the response to the expected final output through conditional keywords, not by API call matching.
Figure 6: Example of the function signature in level-3.
Mechanism M3ToolEval API-Bank level-3
Avg Turns Correct Output Words Avg Turns Correct Output Words
ReAct 8.2 38.1 % 1.86 k 9.5 8.2 % 1.66 k
CodeAct 7.0 49.4 % 1.91 k 8.9 19.2 % 1.82 k
Tree-of-Code (3-3) 1.7 $\downarrow$ 67.1 % $\uparrow$ 0.44 k $\downarrow$ 2.1 $\downarrow$ 38.0 % $\uparrow$ 0.39 k $\downarrow$
Table 1: Performance comparison of the baselines and our ToC in terms of averaged turns, output words, and accuracy on two datasets. Note: all numerical results presented in this paper are rounded.
Baselines. ReAct Yao et al. (2022) combines reasoning and action in a dynamic, step-by-step interaction, providing a flexible approach to task-solving. We use JSON as the action space. CodeAct Wang et al. (2024) utilizes a block of code as the LLM agent’s action, enabling more efficient multi-turn interactions.
Metrics. The evaluation includes accuracy and averaged turns. Accuracy represents the percentage of complex tasks that are correctly solved. We consider the LLM-generated code at the same layer, generated in parallel, as one turn. We also record the average number of output words for the API cost evaluation.
4.2 ToC vs. CodeAct and ReAct
We primarily compare the ToC framework, which is comprised of CodeProgram nodes, with the CodeAct and ReAct framework, which are comprised of steps, using the M3 and the level-3 datasets. For ToC, we randomly sample the LLM and prompt from the LLM list and prompt pool, respectively, at each node exploration. For CodeAct and ReAct, we report the average results across all LLMs used in this paper. Table 1 shows ToC achieves consistent superior performance (nearly 20% higher) with significantly fewer interaction steps and averaged output words (nearly 1/4), highlighting its efficiency in handling complex tool-use scenarios. Specifically, Figure 7 shows the comparison of ReAct, CodeAct, and ToC on the five tasks in the M3, where ToC achieves near-perfect accuracy on all tasks except the web browsing task.
Figure 7: Comparison across five tasks in the M3.
4.2.1 Other multi-turn vs. Our one-turn
Furthermore, we explore the performance of one-layer ToC (1-x) with the fixed model. As a node in ToC, CodeProgram enables the complete solution in a single turn by leveraging code’s ability to handle long logic chains. Table 2 shows that, with a significant advantage in the number of turns (one vs. multi-turn: averaged 7.0/8.9), our performance on some models even surpasses multi-turn CodeAct and ReAct, particularly with the Claude series. Compared to CodeProgram, ie. ToC (1-1), the single layer, three nodes ToC (1-3) with random prompts significantly boosts performance. Its average accuracy already surpasses CodeAct, highlighting the effectiveness of prompt randomness.
4.3 Analysis and Ablation Studies
@@
ReAct CodeAct ToC (1-1) ToC (1-3) ReAct CodeAct ToC (1-1) ToC (1-3)
@@
Table 2: Ablation study of the model exploration. With different fixed models, the detailed performance comparison of ReAct, CodeAct, ablated ToC (1-1) (ie. the CodeProgram node), and ToC (1-3) on the M3ToolEval and API-Bank level-3 datasets is shown. The correctness is reported, with the average number of turns in parentheses.
@@
It seems the "nodes per Layer" contribute more than Layers, likely because our tree structure is designed to enhance exploration. Increasing the number of nodes certainly introduces more diverse prompts and model variations, whereas adding more layers (ie. more turns) mainly accumulates histories without significantly improving decision-making, especially with models that have limited contextual understanding.
Prompt exploration. Ablation results in Table 4 confirm the effectiveness of prompt exploration. By comparing the random model with the fixed model (claude-3-5-sonnet), prompt exploration proves to be more critical in scenarios with lower diversity.
@@
Table 4: Ablation study of the prompt exploration.
@@ 5 Related Work
LLM Code Generation for Complex Tasks. Recent works integrating LLMs with code have largely focused on task completion in programming domains like software development Qian et al. (2024); Wang et al. (2023), programming assistance Islam et al. (2024); Wen et al. (2024b), and scientific problems Chen et al. (2022); Gao et al. (2023); Hong et al. (2024). These studies primarily address pure code generation, where correct task completion only relates to accurate reasoning logic within the code. For example, Chain of Codes Li et al. (2023a) broadens LLM capabilities by enabling "thinking in code." In contrast, our work addresses real-world, zero-shot online complex tasks that involve multiple tool calls. Only CodeAct Wang et al. (2024) treats code as a scalable language to call multiple tools, but their approach is limited by an almost one-turn, one-tool, step-by-step mechanism. This results in stalled thinking and accumulated histories, relying heavily on ground-truth supervision for each step, which is incompatible with zero-shot, online settings. In our framework, every node represents a complete solution that can be directly evaluated via execution supervision without requiring additional labels.
Tree-based Code Generation. A recent work, CodeTree Li et al. (2024b), uses a tree structure to explore the search space of code generation tasks. Unlike our approach, CodeTree focuses on multi-agent searching rather than an end-to-end, self-growing tree. While self-repair trees Olausson et al. (2023) begin with a specification root node, grow into initial programs through separate feedback and repair stages–often bottlenecked by the model’s limited capacity–our approach unifies reasoning (including reflection) and generation in a single cycle at each node, and directly expands the tree with prompt and model exploration. Some contemporaneous works utilizing tree-based search, such as MCTS Xu et al. (2024a); Yu et al. (2024), require multiple rollouts and significant computational resources, making them unsuitable for online, real-time applications. Unlike these methods, our self-growing tree generates multiple valid solutions and directly selects the one closest to the ground truth through a voting mechanism. Additionally, these studies typically focus on tasks with easier-to-obtain process supervision, whereas our work addresses real-world, complex multi-tool datasets.
@@ 6 Conclusion
This paper introduced the Tree-of-Code (ToC) method, which enables self-growing, end-to-end thought-code generation based on successful execution, addressing complex multi-tool online tasks. With efficient model integration and prompt exploration, ToC outperformed baselines on two complex task datasets, improving both efficiency and task-solving performance.
@@
We emphasize that our method operates at the granularity of code "program" rather than "action". However, it is limited in fully open-ended scenarios requiring step-by-step exploration, such as a robot navigating an unfamiliar environment, or in handling tasks with extremely long sequences beyond the capabilities of current reasoning methods, like generating an entire paper. In such cases, it cannot provide a complete final solution. Even though, in practical industrial applications where a predefined toolset is available, CodeProgram’s end-to-end execution remains more efficient for online, zero-shot scenarios, for fewer turns, and for fewer LLM calls.
For larger and more complex system programs in the future, our method may serve as a "subprogram" within the overall solution, similar to a single agent’s role in multi-agent systems.
@@
Acknowledgments
This work was supported by the National Science and Technology Major Project under Grant 2022ZD0116409, and the National Natural Science Foundation of China under Grant 62301559.
@@ B.2.2 NextAction Tool Description
B.3 An example to show the performance of two helpers tools
For web-based queries, ResHandler and NextAction, these two helper tools are essential to ensure that our method can directly generate the final result without requiring intermediate feedback.
To further clarify their effectiveness, we provide a qualitative and specific example demonstrating how they function in practice.
Below is the full code output for the query: "Find the email of Henry Santiago. Answer in the format of ’xxx@xxx.xxx’." from typing import List, Tuple
def find_email(query: str, current_page_content: str, visited_urls: List[str], max_depth: int = 30) -> str:
if max_depth <= 0:
raise Exception("Max search depth reached, email not found.")
# Check if the email is directly available in the current page content
if "Henry Santiago" in current_page_content and "@" in current_page_content:
email = res_handler(f"Extract the email of Henry Santiago from the following text: {current_page_content}")
if email:
return email.strip()
# Identify clickable URLs and navigate to them
next_step_action, reference_page_content = next_action(query, current_page_content, visited_urls)
if next_step_action == "end()":
email = res_handler(f"Extract the email of Henry Santiago from the following text: {reference_page_content}")
if email:
return email.strip()
else:
raise Exception("Email not found in the reference page content.")
elif next_step_action.startswith("click_url"):
url = next_step_action.split(’(’)[1].split(’)’)[0].strip().strip(’"’)
if url not in visited_urls:
visited_urls.append(url)
new_page_content = click_url(url)
return find_email(query, new_page_content, visited_urls, max_depth - 1)
elif next_step_action == "go_to_previous_page()":
new_page_content = go_to_previous_page()
return find_email(query, new_page_content, visited_urls, max_depth - 1)
else:
raise Exception("Unknown next step action encountered.")
# Main function to start the search
def main():
query = "Find the email of Henry Santiago"
initial_page_content = view()
visited_urls = []
try:
email = find_email(query, initial_page_content, visited_urls)
print(email)
except Exception as e:
print(f"Error: {e}")
if __name__ == "__main__":
main()
Functionality of the Helper Tools
For res_handler:
Example usage:
email = res_handler("Extract the email of Henry Santiago from the following text: {current_page_content}")
We provide the content that the final result depends on (i.e., the return value of other functions in the code) as input to this function. It then passes the information to an LLM to generate the final answer.
For next_action:
Example usage:
# Identify clickable URLs and navigate to them
next_step_action, reference_page_content = next_action(query, current_page_content, visited_urls)
This function determines the next action based on: the current page content, the original query, and the list of visited URLs.
Appendix C Visualization of the Table 2